OpenStudy (grace111):

Medal and fan for best answer What is the area of a square with a perimeter of 8cm?

OpenStudy (whpalmer4):

What is the formula for the perimeter of a square? What is the formula for the area of a square?

OpenStudy (whpalmer4):

What is the definition of the perimeter?

OpenStudy (whpalmer4):

When you read a problem, and you don't understand what all of the words mean, your first reaction should be to look them up in your textbook or notes.

OpenStudy (whpalmer4):

I have absolutely no doubt that the text or your instructor defined those terms for you and gave you those formulas, and you aren't expected to derive them on your own.

OpenStudy (grace111):

ok thank you @whpalmer4 for your help that you gave me

OpenStudy (whpalmer4):

Have you found the answer yet?

OpenStudy (grace111):

Im trying to

OpenStudy (whpalmer4):

what's the meaning of the perimeter?

OpenStudy (whpalmer4):

the problem is very easy once we know what the terms and formulas are...

OpenStudy (grace111):

perimeter is when you add all the sides I think

OpenStudy (whpalmer4):

Perimeter: the continuous line forming the boundary of a closed geometric figure If we have a square, we have 4 sides, right? Do we know anything about the lengths of those sides?

OpenStudy (whpalmer4):

Anything interesting about the lengths of the 4 sides of a square?

OpenStudy (whpalmer4):

Doesn't a square have 4 equal sides, by definition?

OpenStudy (grace111):

yes it has

OpenStudy (whpalmer4):

If that's true, what is the length of each side for this square?

OpenStudy (whpalmer4):

that implies that for a square, \[P = 4*s\]where \(P\) is the perimeter, and \(s\) is the length of a side. We have \(P=8\) here, so \[P=4s\]\[8=4s\]\[s=\]

OpenStudy (grace111):

sorry Im bad at maths :( so it takes longer for me to work out word problems.

OpenStudy (whpalmer4):

What number multiplied by 4 gives you 8?

OpenStudy (whpalmer4):

8 / 4 = ?

OpenStudy (grace111):

32

OpenStudy (grace111):

sorry 2

OpenStudy (whpalmer4):

Very good. Now, do you know the formula for the area of a square, if we know the length of the sides? It's the same basic formula as that for the area of a rectangle (because all squares are rectangles, though obviously not all rectangles are squares!): area = length * width for a square, both the length and the width are just the length of any side, so our formula for the area of a square is: \[A = s*s = s^2\] We have \(s = 2 \text{ cm}\) \[A = (2 \text{ cm})*(2 \text{ cm}) =

OpenStudy (whpalmer4):

ah, nuts, that last line should be: \[A = (2 \text { cm})*(2 \text{ cm})=\]

OpenStudy (whpalmer4):

The unit will be \(\text{cm}^2\) (read as "square centimeters")

OpenStudy (whpalmer4):

What's the numeric part of the answer?

OpenStudy (whpalmer4):

\[A = (2 \text{ cm})*(2 \text{ cm}) = 2*2 \text{ cm}^2 = \text{_________} \text{ cm}^2\]

OpenStudy (whpalmer4):

What's the numeric part of the answer?

OpenStudy (whpalmer4):

The unit will be \(\text{cm}^2\) (read as "square centimeters")